The population mean and standard deviation are given below. Find the required probability and determine whether the given sample mean would be considered unusual. For a sample of n=70, the probability of a sample mean being greater than 224 if u=223 and o=5.7 is? Would the given sample mean be considered unusual? The sample mean would/would not be considered unusual because it does/does not lie within a range of usual event.
Answer #1

Let \bar{X} denotes the sample mean.

So, \bar{X}\sim Normal(223, \frac{5.7^2}{70}) or \bar{X}\sim Normal(223, 0.681280^2)

Now,

P(\bar{X}>224) = P(\frac{\bar{X}-223}{0.68128}>\frac{224-223}{0.68128})

\Rightarrow P(\bar{X}>224) =1- P(\frac{\bar{X}-223}{0.68128}\leq \frac{224-223}{0.68128})

\Rightarrow P(\bar{X}>224) =1- P(Z \leq \frac{224-223}{0.68128})

\Rightarrow P(\bar{X}>224) =1- \Phi(\frac{224-223}{0.68128})

\Rightarrow P(\bar{X}>224) =1- \Phi(1.46782)

\Rightarrow P(\bar{X}>224) =1-0.928924 = 0.071076

\Rightarrow P(\bar{X}>224) = 0.071076

Since the probability is not too small, so we fail to conclude that the sample mean can not be greater than 224